OpenStudy (anonymous):

4 grey females, 4 grey males, 2 white females, and 3 white males. As in the book, the biologist selects two mice randomly. (1) What is the probability of selecting two males given that both are grey? (2) What is the probability of selecting 1 mouse of each gender given that both are grey?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Add up all the numbers for the total amount: 4+4+2+3 = 13 Next, figure out the probability of selecting 1 male that is grey, which is 4/13. Doing that a second time would again be 4/13. In order to figure out the probability of randomly selecting two of them and they both are gray males, you multiply the probabilities yielding: 4/13 x 4/13 = 16/169, or about a 9.5% chance.